# Serial and Parallel Two-Column Probing for Mixed-Integer Programming

### v1

@@ 4 Parallelized Two-Column Probing
In this section, we parallelize our two-column probing procedure. We adopt a different approach from that of Gleixner, Gottwald, and Hoen [27] because the two-column probing method involves more global bound updates, aggregations, and implications than one-column probing, and so the communication cost of their scheme may be prohibitive in the extended version. Instead, our method partitions the binary variables and assigns a partition to each thread, which can then be handled independently (Sec. 4.1). Locally, two-column probing (Alg. 4) is applied to variables assigned to each thread. Then the results are reduced/merged globally by using a customized implication analysis procedure (Sec. 4.2).
@@
Chart 3: Two-column Probing (parallel). Orange color rectangles represent new serial subroutines. Gray colors are run in parallel, with each thread applying the algorithms in the dashed rectangle.
@@ 4.1 Variable Partitioning
The main interdependence between the outer for-loop iterations (line 2 of Alg. 4) of the serial implementation occurs in the inner for-loop (line 4). After setting two binary variables to their implied bounds, two-column probing propagates $CT$ to fix additional binary variables. If these $CT$-fixing variables have been probed in previous iterations, two-column probing may fix more variables via $AF$ and $IG$ associated with these $CT$-fixing variables. Thus, for parallelization, we aim to assign variables that conflict with each other to the same thread. Note, however, that maximizing the number of conflicts assigned to each thread is an NP-hard partitioning problem [19, 22]—hence, we propose a greedy heuristic partitioning method described in Alg. 5.
@@ 4.2 Implication Analysis
In one-column probing, a follow-up preprocessing method called implication graph analysis [2] can use the results of probing to update global bounds and detect new aggregations. Implication analysis is similar to one-column probing but drops domain propagation to save time. However, this procedure is notably less effective for serial two-column probing: based on experiments, we observed that implication graph analysis detects very few new aggregations as our two-column probing tends to probe individual variables multiple times, resulting in diminishing returns. However, our parallel two-column algorithm ignores pairs of variables between two threads, and so we adapt implication analysis by running it only on these neglected pairs to try and recover some of the advantage of global probing. After running local two-column probings, we first update global variable bounds by $ub:=\min_{i=1,...,k}\{ub^{i}\}$ and $lb:=\max_{i=1,...,k}\{lb^{i}\}$, where $ub^{i},lb^{i}$ are the upper and lower bounds from the $i$-th thread. We then conduct implication analysis on pairs of variables between threads as described in Alg. 6.
Input: $CT$, $AF$, $IG$, $ub$, $lb$
Output: updated $ub$, $lb$, $AF$
@@
4             Propagate $CT$ for $x_{i}=v$ to fix conflicting variables;
@@ 6             Update $ub$, $lb$, and $AF$;
For each $x_{i},i\in\mathcal{P}$ ($\mathcal{P}$ is defined in line 1 of Alg. 6), after setting it to 00 or $1$, Alg. 6 follows a procedure similar to Alg. 4 but drops domain propagation to save time. In line 4, fixing $x_{i}$ and propagating $CT$ may fix some variables that were assigned to different threads, such as nodes $4$ and $5$ in Fig. 1. Subsequently, propagating $AF$ and $IG$ uses local probing outcomes across threads to detect potential new/stronger implications for $x_{i}$. As a result, probing outcomes from different threads are exchanged and merged into the main thread, restoring some of the global effects that were lost by parallelization.
@@ 4.3 Discussion and Analysis
1 Conduct simple pre-presolve and construct $CT$, $CM$;
@@ 5       Conduct Two-Column Probing for $\{x_{j},j\in\mathcal{I}_{i}\}$;
[…]
@@ 5.1.1 Test Set
Experiments are conducted on the benchmark set of the MIPLIB 2017 Collection [28], consisting of $240$ instances. We exclude $3$ instances due to the memory-handling issues in JuMP. We remove $47$ instances where no pairs of binary variables appears in the same constraint–hence our method terminates with no result. The results below are run on the remaining $190$ instances.
@@ 5.1.2 Software and Hardware
All algorithms are implemented in Julia 1.8.5 [13]. We solve both original MIPs and reduced MIPs (MIPs after applying the two-column probing) SCIP v0.11.14 [4, 3] and an alpha (prototype) version of JuMP 1.28 [36] on an Owens node from the Ohio Supercomputer Center. This node includes 2 sockets, with each socket containing 14-core Intel Xeon E5-2680 CPUs. The node has 128 GB of memory and runs 64-bit Red Hat Enterprise Linux 7.9.0. We use $16$ GB, $16$ cores for the parallel computation, and $1$ core for the SCIP solver.
@@ 5.2 Parameter Setting
[…]
@@ 5.2.1 Measurements
Pre Time measures time spent in two-column probing. This excludes overhead from the reading time of the input file and the time to build and write the reduced model to file by JuMP, but includes building $CM$, $CT$, selecting variable pairs, two-column probing, and the implication analysis. Pairs # recorded how many pairs of binary variables are probed by (parallel) two-column probing.
@@
[…]
We also record the primal-dual gap based on the primal and dual bounds reported by SCIP. To get a gap within a fixed range, here we use the primal-dual gap defined by Berthold [11]. This gap is between $[0,100]$: if SCIP finds an optimal solution or certificates the infeasibility, the gap is 00; if the solver does not find any feasible solution but also cannot certificate the infeasibility, the gap is $100$; otherwise, the gap is calculated as
@@ 5.3 Parallel Performance
111122224444888816161616404040406060606080808080100100100100120120120120140140140140160160160160Number of Threads [$k$]Pairs #Parallel Two-Column ProbingFitting Function: $35k^{\log_{2}(1.45)}$
@@ 5.4 Default SCIP
[…]
@@ 5.5 Two-Column vs One-Column Probing
We also compare two-column probing with one-column probing disabled (setting parameter “propagating/probing/maxuseles” as “false”). Namely experiments from Sec. 5.4 were rerun comparing two-column probing + default SCIP without one-column probing vs. default SCIP. Results are summarized in Table 4.
@@
All 190 103.1% 99.4% 102.3% 103.0% 103.8%
$\geq 10$ 183 104.0% 98.1% 101.4% 103.3% 104.0%
$\geq 100$ 160 96.5% 93.6% 92.5% 96.4% 98.2%
$\geq 1000$ 127 77.0% 66.2% 78.2% 80.9% 80.5%
@@ 6 Conclusion
This paper develops the first two-column probing procedure, with both a serial and parallel version provided. Although our basic implementation does not have the maturity of fully-integrated one-column probing, empirical results with SCIP nonetheless indicate the promise of our approach, especially on more difficult MIPLIB instances. The parallel version appears advantageous compared with our serial method using 8 or 16 threads.

### v2

@@ 4 Parallelized Two-Column Probing
In this section, we parallelize our two-column probing procedure. Our method partitions the binary variables and assigns a partition to each thread, which can then be handled independently (Sec. 4.1). Locally, two-column probing (Alg. 4) is applied to variables assigned to each thread. Then the results are reduced/merged globally by using a customized implication analysis procedure (Sec. 4.2).
@@
Chart 3: Two-column Probing (parallel). Orange color rectangles represent new serial subroutines. White colors are run in parallel, with each thread applying the algorithms in the dashed rectangle.
@@ 4.1 Variable Partitioning
The main interdependence between the outer for-loop iterations (line 2 of Alg. 4) of the serial implementation occurs in the inner for-loop (line 4). After setting two binary variables to their implied bounds, two-column probing propagates $CT$ to fix additional binary variables. If these $CT$-fixing variables have been probed in previous iterations, two-column probing may fix more variables via $AF$ and $IG$ associated with these $CT$-fixing variables. Thus, for parallelization, we aim to assign variables that conflict with each other to the same thread. Note, however, that maximizing the number of conflicts assigned to each thread is an NP-hard partitioning problem [chopra1993partition, devine2006partitioning]—hence, we propose a greedy heuristic partitioning method described in Alg. 5.
@@ 4.2 Implication Analysis
In one-column probing, a follow-up preprocessing method called implication graph analysis [scip] can use the results of probing to update global bounds and detect new aggregations. Implication analysis is similar to one-column probing but drops domain propagation to save time. However, this procedure is notably less effective for serial two-column probing: based on experiments, we observed that implication graph analysis detects very few new aggregations as our two-column probing tends to probe individual variables multiple times, resulting in diminishing returns. However, our parallel two-column algorithm ignores pairs of variables between two threads, and so we adapt implication analysis by running it only on these neglected pairs to try and recover some of the advantage of global probing. After running local two-column probings, we first update global variable bounds by $ub:=\min_{i=1,...,k}\{ub^{i}\}$ and $lb:=\max_{i=1,...,k}\{lb^{i}\}$, where $ub^{i},lb^{i}$ are the upper and lower bounds from the $i$-th thread. We then conduct implication analysis on pairs of variables between threads as described in Alg. 6.
Input: CTCT, AFAF, IGIG, ubub, lblb
Output: updated ubub, lblb, AFAF
@@
4      Propagate CTCT for $x_{i}=v$ to fix conflicting variables;
@@ 6      Update ubub, lblb, and $AF$;
For each $x_{i},i\in\mathcal{P}$ ($\mathcal{P}$ is defined in line 1 of Alg. 6), after setting it to 0 or $1$, Alg. 6 follows a procedure similar to Alg. 4 but drops domain propagation to save time. In line 4, fixing $x_{i}$ and propagating $CT$ may fix some variables that were assigned to different threads, such as nodes $4$ and $5$ in Fig. 1. Subsequently, propagating $AF$ and $IG$ uses local probing outcomes across threads to detect potential new/stronger implications for $x_{i}$. As a result, probing outcomes from different threads are exchanged and merged into the main thread, restoring some of the global effects that were lost by parallelization.
@@ 4.3 Discussion and Analysis
1 Conduct simple pre-presolve and construct CTCT, CMCM;
@@ 5   Conduct Two-Column Probing for $\{x_{j},j\in\mathcal{I}_{i}\}$;
For two-column probing in line 4, we set similar termination conditions discussed in Sec. 3.3.2, but at most probe $\lceil\texttt{Max\_Probe\_Number}/\ln(k+1)\rceil$ pairs of variables, where $k$ is the number of threads. Note that in most cases, the maximum iteration limit and global time limit are relatively loose; thus Alg. 7 can expend similar runtime as the serial version. However, for a given runtime, the parallel version can probe more pairs of variables, as demonstrated by numerical experiments in Sec 5.3. As mentioned at the beginning of the section, local probing is (weakly) less effective than global probing for a given pair. Whether the tradeoff in quantity over quality is worthwhile is ultimately an empirical question, which we explore in the next section.
@@ 5.1.1 Test Set
Experiments are conducted on the benchmark set of the MIPLIB 2017 Collection [miplib], consisting of $240$ instances. We remove 47 instances for which no two-column probing is available—no pair of binary variables appears in the same constraint—as our subroutine simply terminates with minimal time spent and no effect on the solver. The results below are run on the remaining $193/240$ instances.
@@ 5.1.2 Software and Hardware
All algorithms are implemented in Julia 1.8.5 [bezanson2017julia]. We solve both original MIPs and reduced MIPs (MIPs after applying the two-column probing) SCIP v0.11.14 [scip1, scip2] and an alpha (prototype) version of JuMP 1.28 [jump] on an Owens node from the Ohio Supercomputer Center. This node includes 2 sockets, with each socket containing 14-core Intel Xeon E5-2680 CPUs. The node has 128 GB of memory and runs 64-bit Red Hat Enterprise Linux 7.9.0. We use $16$ GB, $16$ cores for the parallel computation, and $1$ core for the SCIP solver.
@@ 5.2 Parameter Setting
Considering our software and hardware, for Alg. 2, we set $\texttt{size\_limit}:=10^{3}$ and $\texttt{work\_limit}:=2\times 10^{8}$; for Alg. 3, $\texttt{Cand\_Number}:=5\times 10^{6}$ and $\texttt{Max\_Probe\_Number}:=10^{3}$; the effort score threshold $\texttt{eff\_threshold}:=1000$ from Sec. 3.3.2.
Instances were predominantly terminated by soft stopping rule. For example, in Tables. 2 and 3, the time limit was hit in either 2/193 or 3/193 instances depending on thread count, with effects on the averages being less than reporting precision ($0.1\%$). Variable limits were only hit in serial implementation (10/193 instances), with similar negligible impact on analysis. The limited impact is because reaching time and variable limits typically coincided with challenging instances that could not be solved irrespective of probing settings.
@@ 5.2.1 Measurements
Pre Time measures time spent in two-column probing. This excludes overhead from the reading time of the input file and the time to build and write the reduced model to file by JuMP, but includes building $CM$, CTCT, selecting variable pairs, two-column probing, and the implication analysis. Pairs # recorded how many pairs of binary variables are probed by (parallel) two-column probing.
@@
$\texttt{Runtime Comp.}:=\frac{\texttt{Pre Time}+\texttt{Rd Time}}{\texttt{Org Time}}.$ (10)
We also record the primal-dual gap based on the primal and dual bounds reported by SCIP. To get a gap within a fixed range, here we use the primal-dual gap defined by Berthold [BERTHOLD2013611]. This gap is between $[0,100]$: if SCIP finds an optimal solution or certificates the infeasibility, the gap is 0; if the solver does not find any feasible solution but also cannot certificate the infeasibility, the gap is $100$; otherwise, the gap is calculated as
@@ 5.3 Parallel Performance
112244881616404060608080100100120120140140160160Number of Threads [$k$]Pairs #Parallel Two-Column ProbingFitting Function: $35k^{\log_{2}(1.45)}$
@@ 5.4 Default SCIP
We employed two-column probing with varying numbers of threads (1, 2, 4, 8, 16) to preprocess 193 instances and generate reduced models. Subsequently, we solved both the original and reduced models using SCIP with default settings, a single thread, a memory limit of 16 GB, and a time limit of $3600-\texttt{Pre Time}$ seconds, where for the original model, time limit is $3600$ seconds. The results of SCIP’s solving process are summarized in Table 2.
@@ 5.5 Two-Column vs One-Column Probing
We also compare two-column probing with one-column probing disabled (setting parameter “propagating/probing/maxuseles” as “false”). Namely experiments from Sec. 5.4 were rerun comparing two-column probing + SCIP without one-column probing vs. default SCIP (including one-column probing). Results are summarized in Table 4.
@@
All 193 103.1% 99.4% 102.3% 103.0% 103.8%
$\geq 10$ 186 104.0% 98.1% 101.4% 103.3% 104.0%
$\geq 100$ 162 96.5% 93.6% 92.5% 96.4% 98.2%
$\geq 1000$ 129 77.0% 66.2% 78.2% 80.9% 80.5%
5.6 Random Seeds with 16 Threads
As a robustness check, we ran additional additional experiments using 5 random seeds per instance to help mitigate solver performance variability [lodi2013performance]. For practical purposes we set a lower total solve time limit of 1800 seconds on these runs, and considered the following configurations:
  • Org: default SCIP (one-column probing enabled).
  • Org-1: SCIP with one-column probing disabled.
  • Two Probing: default SCIP with two-column probing.
  • Two Probing-1: SCIP with two-column probing but one-column probing disabled.
Each percentage in Table 6 is the runtime (gap for Table 7) comparison between the method listed in the column header and the method listed in the row header. For example, the $94.3\%$ in position (1,3) of Table 6 is calculated by $\texttt{Two Probing Runtime}/\texttt{Org Runtime}$, i.e. SCIP + (16-threaded) two-column probing is $6.7\%$ faster than default SCIP across $193$ instances $\times 5$ runs. This is in line with the $95.4\%$ single seed result reported in Table 2, indicating the effectiveness of two-column probing combined with one-column probing. Likewise, the gap comparison of $96.7\%$ in position (1,3) of Table 7 compares favorably to the single seed result of $98.2\%$. Hence, the results of Section 5.4 are supported by these additional experiments. Likewise, results here are consistent with Section 5.6; for example, indicating that two-column probing without one-column probing is slightly less effective than vice versa.
Note that, as with previous subsections, runtime comparison only includes instances that can be solved within the time limit by at least one of the compared methods, while the gap comparison includes instances that cannot be solved within the time limit by at least one method.
Table 6: Runtime Comp. for within time-limit cases between four methods.
Runtime Comp. Org Org-1 Two Probing Two Probing-1
Org - 107% 94.3% 100.6%
Org-1 93.5% - 92.2% 94.1%
Two Probing 106.1% 108.4% - 103.8%
Two Probing-1 99.4% 106.3% 96.4% -
Table 7: Gap Comp. for exceeding time-limit cases between four methods.
Gap Comp. Org Org-1 Two Probing Two Probing-1
Org - 116.7% 96.7% 115.0%
Org-1 85.7% - 85.0% 99.1%
Two Probing 103.4% 117.7% - 117.1%
Two Probing-1 86.9% 100.9% 85.4% -
5.6.1 Per-Instance Comparison of Two-Column vs One-Column
Figure 3 plots instances by runtime comparisons of Two Probing-1/Org-1 on the vertical axis (lower values indicate more time reduction due to two-column probing) and Org/Org-1 on the horizontal axis (lower values indicate more time reduction due to one-column probing). Each point represents a MIPLIB instance (geometric average over 5 random seed runs), with color indicating the number of two-column pairs probed. The figure indicates that when many pairs are probed, the effectiveness of two-column probing on instances is correlated with the effectiveness of one-column probing; otherwise, in instances where we probe fewer pairs, it seems to coincide with minimal one-column probing, hence the cluster of light-coloured points around $100\%$ on the horizontal axis (no one-column effect). This suggests that problem instances that when single-column probing is active, the effects on solve-time would be amplified with the addition of two-column probing. Moreover there are a large number of instances where modest levels of two-column probing is having a substantial effect, but where one-column probing does not.
Figure 3: A scatterplot of total runtime effects: vertical axis is two-column probing, horizontal axis is one-column probing. Each point represents five runs of an instance; shade indicates pairs of variables probed.
5.6.2 Two-Column vs Pairs Probed
Figure 4 plots instances by runtime comparisons of Two Probing/Org, i.e. improvement over default SCIP, on the vertical axis (lower values indicate more time reduction from probing) and the number of pairs probed. Each point represents a MIPLIB instance (geometric average over 5 random seed runs), with color indicating time to solve (the longest time to solve among Two Probing and Org configurations). There does not seem to be a clear pattern, although our previous experiments indicate average overall benefits with higher thread-counts and more pairs probed (see Tables 2 and  3).
Figure 4: A scatterplot of runtime improvements from two-column probing vs total pairs probed. Each point represents five runs of an instance; shade indicates SCIP total time to solve.
@@ 6 Conclusion
This paper develops a two-column probing procedure, with both a serial and parallel version provided. Although our basic implementation does not have the maturity of fully-integrated one-column probing, empirical results with SCIP nonetheless indicate the promise of our approach, especially on more difficult MIPLIB instances. The parallel version appears advantageous compared with our serial method using 8 or 16 threads.
@@
Acknowledgements
This work was funded by the Office of Naval Research under grant N00014-23-1-2632.